# Robustness of Dynamic Quantum Control: Differential Sensitivity Bounds

### v1

@@ III.2 Fidelity Error and Differential Sensitivity
In principle this can be done by numerically evaluating Eq. (16) for a given uncertainty structure $\hat{H}_{\mu}$ and a range of strengths $\delta$ to determine the $\bar{\delta}$ at which $\tilde{\varepsilon}_{\mu}(\bar{\delta})\geq\epsilon$, similar to what is explored in Ref. Khalid et al., 2023a. A drawback of this approach is the computational cost involved, potentially requiring many fidelity evaluations, and the difficulty establishing analytic results.
Here we use the differential sensitivity of the perturbed fidelity error to establish performance guarantees. It was shown in Ref. O’Neil et al., 2023b that we can write
@@
Eq. (18) facilitates two interpretations of the differential sensitivity. If the uncertainty is constant for the entire evolution then $\mathbf{s}_{\mu}$ is a constant vector and $\mathbf{\Gamma}$ can be viewed as a matrix operator that accepts as input a static uncertainty structure, $\mathbf{s_{\mu}}$, and provides as output the sensitivity in that direction. In this case, a static-uncertainty maximum bound on the differential sensitivity at $\delta=0$ can be derived O’Neil et al. (2023b).
If the uncertainty is constant for each time step but varies between time steps then $\hat{H}_{\mu}^{(k)}=\sum_{m=0}^{M}s_{m}^{(k)}\hat{H}_{m}$ with corresponding $\mathbf{s}_{\mu}^{(k)}$ for each time step $k$, and the differential sensitivity can be written as
@@
where $\zeta_{\{\mu\}}$ indicates that $H_{\mu}$ is not fixed but given by the sequence $\{H_{\mu}^{(k)}:k=0,\ldots,\kappa-1\}$. In this case an alternative variable-uncertainty maximum bound, $\mathbf{B_{vu}}$, on the differential sensitivity at $\delta=0$ can be derived O’Neil et al. (2023b):
@@
If the uncertainty can vary from one time step to the next, and is described by the sequence of structure matrices $\set{\hat{H}_{\mu}^{(k)}}_{k=0}^{\kappa-1}$, the variable-uncertainty upper bound on $\|\zeta_{\set{\mu}}\|$ is given $\mathbf{B_{vu}}=\|\set{\bar{\varsigma}^{(k)}}\|_{\ell^{1}}$, the $\ell^{1}$ norm of the sequence $\set{\bar{\varsigma}^{(k)}}$, with $\bar{\varsigma}^{(k)}=\|\mathbf{Z}^{(k)}\|_{2}$. The sequence $\set{\bar{\mathbf{s}}_{\mu}^{(k)}}$ that achieves the bound $\mathbf{B_{vu}}$ is $\bar{\mathbf{s}}_{\mu}^{(k)}={\mathbf{Z}^{(k)}}^{T}/\bar{\varsigma}^{(k)}$ for $0\leq k<\kappa$.
The basic idea behind this proposition is that for each time step we want the size of the differential sensitivity in the direction that maximizes the sensitivity, which is given by the magnitude (or usual $2$-norm) of the vector $\mathbf{Z}^{(k)}$. This yields a sequence of positive numbers over $\kappa$ time steps, and the $\ell_{1}$ norm of this sequence provides an upper bound $\mathbf{B_{vu}}$ for the differential sensitivity. Using $\mathbf{B_{vu}}$ and the associated sequence of maximum sensitivity directions $\set{\bar{\mathbf{s}}_{\mu}^{(k)}}$ allows a directed search to find the maximum perturbation strength $\bar{\delta}$ that guarantees the performance criterion $\tilde{\varepsilon}_{\mu}(\delta)<\epsilon$ is satisfied in any direction provided $\delta\leq\bar{\delta}$ O’Neil et al. (2023b).
Specifically, we quantize the uncertainty size $\delta$ into uniform steps of a given magnitude $\mathbf{d}$. We initialize the sequence of maximum sensitivity directions at $\delta=0$ as $\set{\mathbf{s}_{\mu}^{(k)}(0)}=\set{\bar{\mathbf{s}}_{\mu}^{(k)}}$ following Proposition 1. We then update the perturbed Hamiltonian in this direction at the strength $\mathbf{d}$ as
$\tilde{H}^{(k)}(\delta_{1})=H_{0}+\sum_{m=1}^{M}H_{m}f_{m}^{(k)}+\mathbf{d}\sum_{m=0}^{M}H_{m}f_{m}^{(k)}\alpha_{m}^{(k)}s_{m}^{(k)}(0)$
for each $k$. At each successive step of $n\mathbf{d}=\delta_{n}$ we compute the directions of maximum sensitivity of the fidelity error as $\set{\mathbf{s}^{(k)}_{\mu}(n)}$ for each $k$ as per Proposition 1 where $\mathbf{s}_{\mu}^{(k)}(n)=\left[s_{0}^{(k)}(n),s_{1}^{(k)}(n),\ldots,s_{M}^{(k)}(n)\right]^{T}$ for each $k$ generated by the Hamiltonian $\tilde{H}^{(k)}(\delta_{n})$ at strength $\delta_{n}$. We then continue to iterate on this process until finding the $\delta_{n}$ such that $\tilde{\varepsilon}(\delta_{n})>\epsilon$. The procedure is summarized in Algorithm 1.
@@
3:$\tilde{H}^{(k)}(\delta_{1})=H_{0}+\sum\limits_{m=1}^{M}H_{m}f_{m}^{(k)}+\mathbf{d}\sum\limits_{m=0}^{M}\alpha_{m}^{(k)}\hat{H}_{m}s_{m}^{(k)}(0)\ \forall k$
@@
8:     $\tilde{H}^{(k)}(\delta_{n})=\tilde{H}^{(k)}(\delta_{n-1})+\mathbf{d}\sum\limits_{m=0}^{M}\alpha_{m}^{(k)}\hat{H}_{m}s_{m}^{(k)}(n-1)\ \forall k$
@@ IV Application to Robustness of Quantum Gate Implementation
Table 2: Results of hypothesis tests for correlation between $\mathbf{B_{vu}}$ and the minimum performance-violating perturbation $\bar{\delta}$ based on the Pearson correlation coefficient $r$. ##\## is the number of samples included in the test (i.e., the number of controllers found with $\mathcal{F}>0.99$), $Z_{r}$ is the normalized test statistic, and $p_{r}$ is the $p$-value of the test. Note the strong negative trend between $\mathbf{B_{vu}}$ and $\bar{\delta}$. The ten cases that fail to meet the significance level of $p<0.05$ are shaded.
@@
Following the procedure in Algorithm 1, we can determine the minimum performance-violating perturbation for our test problems, and test the correlation between the local sensitivity at $\delta=0$, as determined by $\mathbf{B_{vu}}$, and the $\bar{\delta}$ that guarantees a performance threshold of $\tilde{\varepsilon}(\delta)<\epsilon$ by a one-tailed hypothesis test. We test for negative correlation indicative of a trade-off in local sensitivity around $\delta=0$ and the maximum allowable perturbation $\bar{\delta}$ that violates the performance criteria at larger perturbation strength. Using the Pearson $r$ Illowsky and Dean (2022) as the correlation measure reveals a strong negative correlation between $\mathbf{B_{vu}}$ and $\bar{\delta}$. Establishing the null hypothesis $\mathcal{H}_{0}$ as no trend between $\mathbf{B_{vu}}$ and $\bar{\delta}$, and $\mathcal{H}_{1}$ as a negative correlation between the metrics with a statistical significance level of $95\%$ ($p=0.05)$, Table 2 shows rejection of $\mathcal{H}_{0}$ in favor of $\mathcal{H}_{1}$ for the majority of test cases. Of the $43$ test cases involving quasi-newton-optimized controllers, all show a negative correlation. Ten tests fail to meet the statistical significance criteria, of which eight are Heisenberg-coupled chains while the other two are Ising chains. While not an exact analytic relation, this suggests that the strictly local differential sensitivity bound $\mathbf{B_{vu}}$ is statistically correlated with robustness to larger perturbation values, agnostic of the structure or direction of the perturbation. Specifically, a higher value of $\mathbf{B_{vu}}$, indicating a greater potential differential sensitivity, correlates with a smaller value of the minimum performance-violating perturbation $\bar{\delta}$, indicative of a smaller margin for acceptable uncertainty.
Having established an upper bound $\mathbf{B_{vu}}$ on the differential sensitivity of the fidelity error and demonstrated its utility as a local measure of sensitivity and indicator of robustness, we examine the correlation between robustness quantified by $\mathbf{B_{vu}}$ and performance quantified by the nominal fidelity error $\varepsilon$ for the control problems with different characteristics introduced earlier.
@@ IV.1 Fidelity Error versus Differential Sensitivity Trends
(a) $\mathbf{B_{vu}}$ versus $\varepsilon$ for Problem $4$ controllers with different gate operation times $t_{f}$ and time resolutions $\kappa$.
(b) $\mathbf{B_{vu}}$ versus $\varepsilon$ for Problem $9$ controllers with different gate operation times $t_{f}$ and time resolutions $\kappa$
@@
Figure 1: Plots of $\mathbf{B_{vu}}$ versus $\varepsilon$ on a log-log scale for the controllers associated with Problem $4$ and Problem $9$ show a strong positive trend between the differential sensitivity bound and fidelity error for a range of gate operation times and time-resolutions.
The upper bound on the differential sensitivity $\mathbf{B_{vu}}$ and the nominal fidelity error $\varepsilon$ show a strong positive correlation. In particular, measuring the linear correlation between these metrics by the Pearson $r$ and testing for positive correlation via a one-tailed hypothesis test indicates this concordant relationship. Of the $86$ cases considered, only one does not meet the significance level of $95\%$. For the remaining $85$, the null-hypothesis $\mathcal{H}_{0}$ of no correlation is rejected in favor of $\mathcal{H}_{1}$, i.e., positive correlation between $\mathbf{B_{vu}}$ and $\varepsilon$. This trend is readily evident in Figs 1a and 1, which display $\mathbf{B_{vu}}$ as function of $\varepsilon$ on a log-log scale for the different $(t_{f},\kappa)$ cases considered for Problems $4$ and $9$, respectively. These plots are characteristic of the results for all problems. In addition to the highly linear trend observed, we note a clustering of controllers in the lower-left corner of the plot in Fig. 1a, the space occupied by controllers with good performance (low error) and low sensitivity. Furthermore, we find that this cluster is dominated by long-time and high-$\kappa$ controllers, indicating that long control time and higher time resolution of the controls leads to more robust controllers.
@@
These findings are significant in that a considerable amount of work has been done on trying to simultaneously optimize performance and robustness Kosut et al. (2013); Koswara et al. (2021a, b); Khaneja et al. (2005b); Ge et al. (2020). Much of this work is predicated on the assumption that there is competition between performance (low fidelity error) and robustness of the controllers. The concordant relationship we observe here, however, suggests that the best performing controllers are also the most robust, at least with regard to the chosen differential sensitivity robustness measure for the problems considered. This concordance suggests that it may not be necessary to perform computationally expensive average fidelity optimization in many cases. Furthermore, this robustness is not limited to uncertainty in the controls but extends to uncertainty in the system and interaction Hamiltonians and other structured perturbations. This positive correlation between the fidelity error and the differential sensitivity has been seen before in the context of time-invariant control of quantum state transfer Jonckheere et al. (2017); Schirmer et al. (2018b). One limitation is the applicability of the differential sensitivity bound, which by its nature is mainly valid in the small perturbation regime. For larger perturbations, it is likely that trade-offs between performance and robustness will emerge. However, it has also been noted that broader peaks in the optimization landscape may be both more robust with regard to noise, and easier to find, especially for certain algorithms such as reinforcement learning Khalid et al. (2021, 2023b), which suggests that, at least in some cases, robust solutions with regard to noise may actually be easier to find than less robust solutions. Less intuitively, it appears that robustness with regard to control imperfections often also implies robustness with regard to other perturbations. An interesting question here is whether there is a maximum uncertainty for a problem where the fidelity error, sensitivity and other robustness measures such as the RIM agree, and how problem-specific such as maximum would be.
@@ IV.2 Sensitivity and System Size
To examine the relation between the nominal fidelity error and the differential sensitivity bound as a function of system size, we compare the results for Problems $1$ through $4$, which use the same system type and control implementation (fixed Ising coupling with individual spin addressability) for system sizes increasing from two to five qubits. Fig. 2 shows the relationship between $\mathbf{B_{vu}}$ and $\varepsilon$ for Problems $1$ to $4$ with controllers optimized for two common values of $\Delta=t_{f}/\kappa$. Fig. 2a shows the results for controllers with $\Delta=t_{f}/\kappa\approx 0.1$ ($\Delta\in[0.094,0.117]$). Fig. 2b shows $\mathbf{B_{vu}}$ versus $\varepsilon$ for controllers with $t_{f}/\kappa\in[0.195,0.234]$. In both cases the larger system (the five-qubit chain of Problem $4$) yields controllers that exhibit both higher fidelity error and greater differential sensitivity as measured by $\mathbf{B_{vu}}$. At the other end of the spectrum, Fig. 2a shows the controllers optimized for Problem $1$ (the less-complex two-qubit system) clustering in the lower-left, the region for the lowest error and differential sensitivity bounds. Though the basic trend of controllers for larger systems showing a less desirable performance-robustness profile is seen in both figures, the difference is less pronounced for the larger values of $\Delta$. Interestingly, in Fig. 2a we observe a cluster of Problem $4$ controllers with similar profiles to those of Problem $2$, and in Fig. 2b we see the same similarity for Problem $3$ and Problem $4$ controllers. This shows that while a more complex (larger) system generally results in lower fidelity and higher sensitivity controllers for a similar control architecture, it is not the only determining factor. Rather a combination of factors, e.g., $t_{f}$ and $\kappa$, also affect controller robustness, which can be leveraged to achieve the best results.
(a) $\mathbf{B_{vu}}$ versus $\varepsilon$ on a log-log scale for controllers with $t_{f}/\kappa\approx 0.1$ for Problems 11112222 and $4$.
(b) $\mathbf{B_{vu}}$ versus $\varepsilon$ on a log-log scale for the controllers with $t_{f}/\kappa\approx 0.2$ in Problems $2$ to $4$
@@
Figure 2: Plots of $\mathbf{B_{vu}}$ versus $\varepsilon$ on a log-log scale for controllers with similar $\Delta=t_{f}/\kappa$. For increasing system size and a common value of $t_{f}/\kappa$, robustness decreases with system size. This observation is more pronounced for the $t_{f}/\kappa\approx 0.1$ case of the top panel than for $t_{f}/\kappa\approx 0.2$ case of the lower panel.
(a) Plot of $\mathbf{B_{vu}}$ versus $\varepsilon$ on a log-log scale for the controllers in Problems $6$ and $9$ with $t_{f}/\kappa<0.1750$
(b) Plot of $\mathbf{B_{vu}}$ versus $\varepsilon$ on a log-log scale for the controllers in Problems $6$ and $9$ with $t_{f}/\kappa>0.200$.
@@
Figure 3: Plot of $\mathbf{B_{vu}}$ versus $\varepsilon$ on a log-log scale for controllers for Problem $6$ and $9$ grouped by similar values of $t_{f}/\kappa$. The top plot shows that individual spin addressability yields controllers with a superior robustness-performance profile than those for control on only the first spin for $t_{f}/\kappa$ between $0.1562$ and $0.1750$. However, the bottom plot shows that for $\kappa$ in the range of $0.200$ to $0.2344$, both control implementations yield similar performance-robustness profiles.
(a) Plot of $\mathbf{B_{vu}}$ versus $\varepsilon$ on a log-log scale for the controllers in Problems $5$ and $8$ with $t_{f}/\kappa<0.1750$
(b) Plot of $\mathbf{B_{vu}}$ versus $\varepsilon$ on a log-log scale for the controllers in Problems $5$ and $8$ with $t_{f}/\kappa>0.200$.
@@
Figure 4: Plot of $\mathbf{B_{vu}}$ versus $\varepsilon$ on a log-log scale for controllers from Problems $5$ and $8$ grouped by similar values of $t_{f}/\kappa$. In the top plot, we see that the implementation of individual qubit control yields controllers with a superior robustness-performance profile than those with control on only the first qubit for $t_{f}/\kappa$ between $0.1562$ and $0.1750$. However, the bottom plot shows that for $\kappa$ in the range of $0.200$ to $0.2344$, both control implementations yield similar performance, with only slightly smaller differential sensitivity for the individual spin addressable case.
(a) Controllers for five-qubit gate Problems $4$ and $7$ for a range of $t_{f}$ and $\kappa$ combinations.
(b) Problem $7$ controllers for $\kappa=1000$ and increasing $t_{f}$.
(c) Problem $4$ controllers for fixed $\kappa=128$ and increasing $t_{f}$.
@@
Figure 5: Comparison of $\mathbf{B_{vu}}$ versus $\varepsilon(t_{f})$ for the Ising ZZ $5$-Chains of Problem 4 and Problem 7. While problem problems have the same target QFT gate, Problem 4 implements individual spin addressable control, and Problem 7 implements simultaneous control on all spins.
(a) Plot of $\mathbf{B_{vu}}$ versus $\varepsilon$ on a log-log scale for the controllers in Problems 22225555, and $6$ with $t_{f}=7$
(b) Plot of $\mathbf{B_{vu}}$ versus $\varepsilon$ on a log-log scale for the controllers in Problems $8$ and $9$ with $t_{f}=15$.
@@
Figure 6: Plot of $\mathbf{B_{vu}}$ versus $\varepsilon$ for three-chain controllers grouped by similar control implementation. The top plot shows that for individual spin-addressability, the Ising-coupled system results in greater fidelity error than the Heisenberg coupled system. Additionally, despite the difference in target gate between Problem $5$ and Problem $6$, both yield controllers with comparable performance-robustness profiles. The bottom-plot shows the results for Problems $8$ and $9$ where the $t_{f}/\kappa$ ratio impacts the performance and robustness more than the difference in target gate.

### v2

@@ III.2 Fidelity Error and Differential Sensitivity
In principle this can be done by numerically evaluating Eq. (15) for a given uncertainty structure ${H}_{\mu}$ and a range of strengths $\delta$ to determine the $\bar{\delta}$ at which $\tilde{\varepsilon}_{\mu}(\bar{\delta})\geq\epsilon$, similar to what is explored in Ref. 16. A drawback of this approach is the computational cost involved, potentially requiring many fidelity evaluations, and the difficulty establishing analytic results.
Here we use the differential sensitivity of the perturbed fidelity error to establish performance guarantees. If the uncertainty structure is constant over time, i.e., $\mathbf{s}^{(k)}_{\mu}=\mathbf{s}_{\mu}$ for all $k$, then Ref. 26 shows that we can write
@@
Eq. (17) facilitates two interpretations of the differential sensitivity. If the uncertainty is constant for the entire evolution then $\mathbf{s}_{\mu}$ is a constant vector and $\mathbf{\Gamma}$ can be viewed as a matrix operator that accepts as input a static uncertainty structure, $\mathbf{s_{\mu}}$, and provides as output the sensitivity in that direction. In this case, a static-uncertainty maximum bound on the differential sensitivity at $\delta=0$ can be derived [26].
If the uncertainty is constant for each time step but varies between time steps then we have the more general case $H_{\mu}^{(k)}=\sum_{m=0}^{M}s_{m}^{(k)}\hat{H}_{m}\alpha_{m}^{(k)}$ with corresponding $\mathbf{s}_{\mu}^{(k)}$ for each time step $k$. The differential sensitivity can thus be written as
@@
where $\zeta_{\{\mu\}}$ indicates that $H_{\mu}$ is not fixed but given by the sequence $\{H_{\mu}^{(k)}:k=0,\ldots,\kappa-1\}$. In this case an alternative variable-uncertainty (vu) maximum bound, $\mathcal{B_{\text{vu}}}$, on the differential sensitivity at $\delta=0$ can be derived [26]:
@@
If the uncertainty can vary from one time step to the next, and is described by the sequence of structure matrices $\set{\hat{H}_{\mu}^{(k)}}_{k=0}^{\kappa-1}$, the variable-uncertainty upper bound on $\|\zeta_{\set{\mu}}\|$ is given $\mathcal{B_{\text{vu}}}=\|\set{\bar{\varsigma}^{(k)}}\|_{\ell^{1}}$, the $\ell^{1}$ norm of the sequence $\set{\bar{\varsigma}^{(k)}}$, with $\bar{\varsigma}^{(k)}=\|\mathbf{Z}^{(k)}\|_{2}$. The sequence $\set{\bar{\mathbf{s}}_{\mu}^{(k)}}$ that achieves the bound $\mathcal{B_{\text{vu}}}$ is $\bar{\mathbf{s}}_{\mu}^{(k)}={\mathbf{Z}^{(k)}}^{T}/\bar{\varsigma}^{(k)}$ for $0\leq k<\kappa$.
The basic idea behind this proposition is that for each time step we want the size of the differential sensitivity in the direction that maximizes the sensitivity, which is given by the magnitude (or usual $2$-norm) of the vector $\mathbf{Z}^{(k)}$. This yields a sequence of positive numbers over $\kappa$ time steps, and the $\ell_{1}$ norm of this sequence provides an upper bound $\mathcal{B}_{\text{vu}}$ for the differential sensitivity. Using $\mathcal{B}_{\text{vu}}$ and the associated sequence of maximum sensitivity directions $\set{\bar{\mathbf{s}}_{\mu}^{(k)}}$ allows a directed search to find the maximum perturbation strength $\bar{\delta}$ that guarantees the performance criterion $\tilde{\varepsilon}_{\mu}(\delta)<\epsilon$ is satisfied in any direction provided $\delta\leq\bar{\delta}$ [26].
Specifically, we quantize the uncertainty size $\delta$ into uniform steps of a given magnitude $\mathtt{d}$. We choose the step size $\mathtt{d}$ small enough so that evaluation of the fidelity for a perturbation of size $\mathtt{d}$ in the worst-case direction $\set{\mathbf{s}^{(k)}_{\mu}}$ remains small. Specifically, we define small as the relative error between the nominal fidelity error $\varepsilon$ and the perturbed fidelity error as $\tilde{\varepsilon}_{\mu}(\mathtt{d})<1/10$. We initialize the sequence of maximum sensitivity directions at $\delta=0$ as $\set{\mathbf{s}_{\mu}^{(k)}(0)}=\set{\bar{\mathbf{s}}_{\mu}^{(k)}}$ following Proposition 1. We then update the perturbed Hamiltonian in this direction at the strength $\mathtt{d}$ as
$\tilde{H}^{(k)}(\delta_{1})=H_{0}+\sum_{m=1}^{M}H_{m}f_{m}^{(k)}+\mathtt{d}\sum_{m=0}^{M}\hat{H}_{m}f_{m}^{(k)}\alpha_{m}^{(k)}s_{m}^{(k)}(0)$
for each $k$. At each successive step of $n\mathtt{d}=\delta_{n}$ we compute the directions of maximum sensitivity of the fidelity error as $\set{\mathbf{s}^{(k)}_{\mu}(n)}$ for each $k$ as per Proposition 1 where $\mathbf{s}_{\mu}^{(k)}(n)=\left[s_{0}^{(k)}(n),s_{1}^{(k)}(n),\ldots,s_{M}^{(k)}(n)\right]^{T}$ for each $k$ generated by the Hamiltonian $\tilde{H}^{(k)}(\delta_{n})$ at strength $\delta_{n}$. We then continue to iterate on this process until finding the $\delta_{n}$ such that $\tilde{\varepsilon}_{\mu}(\delta_{n})>\epsilon$. The procedure is summarized in Algorithm 1.
@@
3:$\tilde{H}^{(k)}(\delta_{1})=H_{0}+\sum\limits_{m=1}^{M}H_{m}f_{m}^{(k)}+\mathtt{d}\sum\limits_{m=0}^{M}\alpha_{m}^{(k)}\hat{H}_{m}s_{m}^{(k)}(0)\ \forall k$
@@
8:     $\tilde{H}^{(k)}(\delta_{n})=\tilde{H}^{(k)}(\delta_{n-1})+\mathtt{d}\sum\limits_{m=0}^{M}\alpha_{m}^{(k)}\hat{H}_{m}s_{m}^{(k)}(n-1)\ \forall k$
@@ IV Application to Robustness of Quantum Gate Implementation
Table 2: Results of hypothesis tests for correlation between $\mathcal{B}_{\text{vu}}$ and the minimum performance-violating perturbation $\bar{\delta}$ based on the Pearson correlation coefficient $r$. ##\## is the number of samples included in the test (i.e., the number of controllers found with $\mathcal{F}>0.99$), $Z_{r}$ is the normalized test statistic, and $p_{r}$ is the $p$-value of the test. Note the strong negative trend between $\mathcal{B}_{\text{vu}}$ and $\bar{\delta}$. The ten cases that fail to meet the significance level of $p<0.05$ are shaded.
@@
Following the procedure in Algorithm 1, we can determine the minimum performance-violating perturbation for our test problems, and test the correlation between the local sensitivity at $\delta=0$, as determined by $\mathcal{B}_{\text{vu}}$, and the $\bar{\delta}$ that guarantees a performance threshold of $\tilde{\varepsilon}(\delta)<\epsilon$ by a one-tailed hypothesis test. We test for negative correlation indicative of a trade-off in local sensitivity around $\delta=0$ and the maximum allowable perturbation $\bar{\delta}$ that violates the performance criteria at larger perturbation strength. Using the Pearson $r$ [44] as the correlation measure reveals a strong negative correlation between $\mathcal{B}_{\text{vu}}$ and $\bar{\delta}$. Establishing the null hypothesis $\mathcal{H}_{0}$ as no trend between $\mathcal{B}_{\text{vu}}$ and $\bar{\delta}$, and $\mathcal{H}_{1}$ as a negative correlation between the metrics with a statistical significance level of $95\%$ ($p=0.05)$, Table 2 shows rejection of $\mathcal{H}_{0}$ in favor of $\mathcal{H}_{1}$ for the majority of test cases. Of the $43$ test cases involving quasi-newton-optimized controllers, all show a negative correlation. Ten tests fail to meet the statistical significance criteria, of which eight are Heisenberg-coupled chains while the other two are Ising chains. While not an exact analytic relation, this suggests that the strictly local differential sensitivity bound $\mathcal{B}_{\text{vu}}$ is statistically correlated with robustness to larger perturbation values, agnostic of the structure or direction of the perturbation. Specifically, a higher value of $\mathcal{B}_{\text{vu}}$, indicating a greater potential differential sensitivity, correlates with a smaller value of the minimum performance-violating perturbation $\bar{\delta}$, indicative of a smaller margin for acceptable uncertainty.
Having established an upper bound $\mathcal{B}_{\text{vu}}$ on the differential sensitivity of the fidelity error and demonstrated its utility as a local measure of sensitivity and indicator of robustness, we examine the correlation between robustness quantified by $\mathcal{B}_{\text{vu}}$ and performance quantified by the nominal fidelity error $\varepsilon$ for the control problems with different characteristics introduced earlier.
@@ IV.1 Fidelity Error versus Differential Sensitivity Trends
Figure 1: Plots of $\mathcal{B}_{vu}$ versus $\varepsilon$ on a log-log scale for the controllers associated with (a) Problem $4$ and (b) Problem $9$ with different gate operation times $t_{f}$ and time resolutions $\kappa$. These plots show a strong positive trend between the differential sensitivity bound and fidelity error for a range of gate operation times and time-resolutions.
@@
These findings are significant in that a considerable amount of work has been done on trying to simultaneously optimize performance and robustness [45, 36, 46, 21, 47]. Much of this work is predicated on the assumption that there is competition between performance (low fidelity error) and robustness of the controllers. The concordant relationship we observe here, however, suggests that the best performing controllers are also the most robust, at least with regard to the chosen differential sensitivity robustness measure for the problems considered. This concordance suggests that it may not be necessary to perform computationally expensive average fidelity optimization in many cases. Furthermore, this robustness is not limited to uncertainty in the controls but extends to uncertainty in the system and interaction Hamiltonians and other structured perturbations. This positive correlation between the fidelity error and the differential sensitivity has been seen before in the context of time-invariant control of quantum state transfer [48, 49]. One limitation is the applicability of the differential sensitivity bound, which by its nature is mainly valid in the small perturbation regime. For larger perturbations, it is likely that trade-offs between performance and robustness will emerge. However, it has also been noted that broader peaks in the optimization landscape may be both more robust with regard to noise, and easier to find, especially for certain algorithms such as reinforcement learning [50, 51], which suggests that, at least in some cases, robust solutions with regard to noise may actually be easier to find than less robust solutions. Less intuitively, it appears that robustness with regard to control imperfections often also implies robustness with regard to other perturbations. An interesting question here is whether there is a maximum uncertainty for a problem where the fidelity error, sensitivity and other robustness measures such as the RIM agree, and how problem-specific such as maximum would be.
@@ IV.2 Sensitivity and System Size
To examine the relation between the nominal fidelity error and the differential sensitivity bound as a function of system size, we vary the number of qubits while holding other variables such as the coupling model and control implementation constant. To this end, we compare the results for Problems $1$ through $4$, which employ the same system type and control implementation (fixed Ising coupling with individual spin addressability) for system sizes increasing from two to five qubits. Fig. 2 shows the relationship between $\mathcal{B}_{\text{vu}}$ and $\varepsilon$ for Problems $1$ to $4$ with controllers optimized for two common values of $\Delta=t_{f}/\kappa$. Fig. 2a shows the results for controllers with $\Delta=t_{f}/\kappa\approx 0.1$ ($\Delta\in[0.094,0.117]$). Fig. 2b shows $\mathcal{B}_{\text{vu}}$ versus $\varepsilon$ for controllers with $t_{f}/\kappa\in[0.195,0.234]$. In both cases the larger system (the five-qubit chain of Problem $4$) yields controllers that exhibit both higher fidelity error and greater differential sensitivity as measured by $\mathcal{B}_{\text{vu}}$. At the other end of the spectrum, Fig. 2a shows the controllers optimized for Problem $1$ (the less-complex two-qubit system) clustering in the lower-left, the region for the lowest error and differential sensitivity bounds. Though the basic trend of controllers for larger systems showing a less desirable performance-robustness profile is seen in both figures, the difference is less pronounced for the larger values of $\Delta$. Interestingly, in Fig. 2a we observe a cluster of Problem $4$ controllers with similar profiles to those of Problem $2$, and in Fig. 2b we see the same similarity for Problem $3$ and Problem $4$ controllers. This shows that while a more complex (larger) system generally results in lower fidelity and higher sensitivity controllers for a similar control architecture, it is not the only determining factor. Rather a combination of factors, e.g., $t_{f}$ and $\kappa$, also affect controller robustness, which can be leveraged to achieve the best results.
@@
Figure 2: Plots of $\mathcal{B}_{\text{vu}}$ versus $\varepsilon$ on a log-log scale for controllers with similar $\Delta=t_{f}/\kappa$ where (a) shows $t_{f}/\kappa\approx 0.1$ for Problems 11112222 and $4$ and (b) shows $t_{f}/\kappa\approx 0.2$ in Problems $2$ to $4$. These plots show that for increasing system size and a common value of $t_{f}/\kappa$, robustness decreases with system size. This observation is more pronounced for the $t_{f}/\kappa\approx 0.1$ case in (a) than for $t_{f}/\kappa\approx 0.2$ case in (b).
@@
Figure 3: Plot of $\mathcal{B}_{\text{vu}}$ versus $\varepsilon$ on a log-log scale for controllers for Problem $6$ and $9$ grouped by similar values of $t_{f}/\kappa$, where (a) shows $t_{f}/\kappa<0.1750$ and (b) shows $t_{f}/\kappa>0.200$. The plot in (a) shows that individual spin addressability yields controllers with a superior robustness-performance profile than those for control on only the first spin for $t_{f}/\kappa$ between $0.1562$ and $0.1750$. However, the plot in (b) shows that for $\kappa$ in the range of $0.200$ to $0.2344$, both control implementations yield similar performance-robustness profiles.
@@
Figure 4: Plot of $\mathcal{B}_{\text{vu}}$ versus $\varepsilon$ on a log-log scale for controllers from Problems $5$ and $8$ for (a) $t_{f}/\kappa<0.1750$ and (b) $t_{f}/\kappa>0.200$. In (a), we see that the implementation of individual qubit control yields controllers with a superior robustness-performance profile than those with control on only the first qubit for $t_{f}/\kappa$ between $0.1562$ and $0.1750$. However, (b) shows that for $\kappa$ in the range of $0.200$ to $0.2344$, both control implementations yield similar performance, with only slightly smaller differential sensitivity for the individual spin addressable case.
@@
Figure 5: Comparison of $\mathcal{B}_{\text{vu}}$ versus $\varepsilon(t_{f})$ for the Ising ZZ $5$-Chains of Problem 4 and Problem 7. (a) Controllers for five-qubit gate Problems $4$ and $7$ for a range of $t_{f}$ and $\kappa$ combinations. (b) Problem $7$ controllers for $\kappa=1000$ and increasing $t_{f}$. (c) Problem $4$ controllers for fixed $\kappa=128$ and increasing $t_{f}$. While both problems have the same target QFT gate, Problem 4 implements individual spin addressable control, and Problem 7 implements simultaneous control on all spins.
@@
Figure 6: Plot of $\mathcal{B}_{\text{vu}}$ versus $\varepsilon$ for three-chain controllers grouped by similar control implementation. (a) Plot of $\mathcal{B}_{\text{vu}}$ versus $\varepsilon$ on a log-log scale for the controllers in Problems 22225555, and $6$ with $t_{f}=7$. (b) Plot of $\mathcal{B}_{\text{vu}}$ versus $\varepsilon$ on a log-log scale for the controllers in Problems $8$ and $9$ with $t_{f}=15$. The plot in (a) shows that for individual spin-addressability, the Ising-coupled system results in greater fidelity error than the Heisenberg coupled system. Additionally, despite the difference in target gate between Problem $5$ and Problem $6$, both yield controllers with comparable performance-robustness profiles. The plot in (b) shows the results for Problems $8$ and $9$ where the $t_{f}/\kappa$ ratio impacts the performance and robustness more than the difference in target gate.